### v1

# On the enumeration of connected sets in finite lattice graphs

Abstract
A connected set in a graph is a nonempty subset of vertices whose induced subgraph is connected. In an infinite lattice, a connected set is often referred to as a lattice animal, whose enumeration up to isomorphism is a classical problem in both combinatorics and statistical physics. In this paper, we focus on the enumeration of connected sets in finite lattice graphs, providing a link between combinatorial counting and structural connectivity in the system.
For any positive integers $m,n$, let $N(P_{m}\times P_{n})$ and $N(C_{m}\times P_{n})$ denote the number of all connected sets in the $(m\times n)$-lattice graph and $(m\times n)$-cylindrical lattice graph, respectively. In 2020, Vince derived enumeration formulas for $N(P_{m}\times P_{2})$ and $N(C_{m}\times P_{2})$, and highlighted the increasing difficulty of extending these calculation results to larger (cylindrical) lattice graphs. Recently, the authors of this paper have developed a method based on multi-step recurrence formulas to obtain enumeration results for $N(P_{m}\times P_{n})$ with $m\leq 4$. In this article, we apply a similar approach to derive the enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$. Further, for the general case, we establish an explicit and tight lower bound on the number of connected sets in any Cartesian product graph with the path of order $n$ as a factor, by employing the transfer matrix method on a subclass of connected sets. Based on this, corresponding asymptotic analysis is provided for some lattice graphs and other graphs of this type as $n$ tends to infinity.
Keywords: lattice animal; lattice; connected set; recurrence formula.
1 Introduction
The enumeration of connected structures on lattices has long attracted interest in both combinatorics and statistical physics [10, 11, 15]. Examples include the dimer problem that counts perfect matchings [13, 23], and the counting of spanning forests and acyclic orientations [7, 1], independent sets [17, 2], and other graph parameters for geometric structures such as the Sierpiski gasket [4, 5, 6, 24]. In particular, lattice animals (finite sets of connected lattice sites or bonds) appear naturally in problems such as percolation, branching processes, and polymer networks as it models connected structures such as clusters and polymer configurations. Most research on the enumeration of lattice animals focuses on distinct animals of a given size on an infinite lattice up to translation and other symmetries, as well as the asymptotic behavior as the size of the animal tends to infinity.
In this paper, we consider a related but complementary problem. Rather than focusing on an infinite lattice and distinct animals, we are interested in the number of all animals within a finite lattice region, which reflects the connectivity of the finite lattice regarding its size and geometry. This research is also motivated by, and originates from, relevant studies on the enumeration of spanning trees in lattices [8, 22]. From a statistical mechanics perspective, enumerating all animals in a finite region is similar to listing all microstates of a discrete system, with each animal corresponding to a distinct microconfiguration. In this way, our study connects combinatorial enumeration with finite-system statistical mechanics, linking structural counts directly to system connectivity and size effects. Furthermore, this approach may offer insights relevant to percolation and other lattice-based models in statistical physics.
In general, for any graph $G=(V(G),E(G))$, a connected set $C$ of $G$ is a subset of $V(G)$ such that the subgraph of $G$ induced by $C$, denoted by $G[C]$, is connected. Clearly, when $G$ is a lattice, any connected set of $G$ is an animal and vice versa. Denote by $\mathcal{C}(G)$ the set of all the connected sets of $G$, let $N(G)=|\mathcal{C}(G)|$, and let $c(G)=\sqrt[|V(G)|]{N(G)}$. Note that $c(G)$ is a parameter that has been independently introduced or employed by many researchers to study the asymptotic behavior of $N(G)$ [3, 9, 12].
In the literature, research on counting connected sets has been extensive and approached from various perspectives. For instance, Kangas et al. [12] determined upper bounds of $N(G)$ for graphs $G$ with the maximum dergee at most 5; Luo et al. [14] established sharp upper and lower bounds of $N(G)$ for any $G$ based on certain graph parameters; Cambie et al. [3] recently studied the maximum number of connected sets for regular graphs; Vince [19] conducted studies on several parameters derived from $N(G)$ and proved that $P_{n}$ minimizes the average order of connected sets over all connected graphs of order $n$; Haslegrave [9] bounded the connected set density away from 1 for graphs with the minimum degree at least 3; and Vince [20] showed that if $G$ is a connected vertex transitive graph, then the average order of the connected sets of $G$ is at least $|V(G)|/2$.
Back to the problem we are concerned with, let $P_{n}$ and $C_{n}$ be the path and cycle of order $n$ for any positive integer $n$, respectively. The Cartesian product of any pair of graphs $G$ and $H$ is denoted by the graph $G\times H$, where $V(G\times H)=V(G)\times V(H)$ and $(x_{1},y_{1})(x_{2},y_{2})\in E(G\times H)$ if and only if either $x_{1}x_{2}\in E(G)$ and $y_{1}=y_{2}$, or $y_{1}y_{2}\in E(H)$ and $x_{1}=x_{2}$. Then $P_{m}\times P_{n}$ and $C_{m}\times P_{n}$ are lattice graphs, and are known as the $(m\times n)$-lattice graph and the $(m\times n)$-cylindrical lattice graph, respectively. See two examples in Figure 1.
[0.35]Refer to caption
Figure 1: $C_{4}\times P_{4}$ and $C_{5}\times P_{4}$
In 2020, Vince [21] derived explicit enumeration formulas for connected sets in $P_{n}\times P_{2}$ and $C_{n}\times P_{2}$ that
$N(P_{n}\times P_{2})=\frac{\beta(n+3)-4n-7}{2}~~\text{and}~~N(C_{n}\times C_{2})=1-3n+2\beta(n)+3n\overline{\beta}(n),$
where $\overline{\beta}(n)$ and $\beta(n)$ are Pell number and Pell-Lucas numbers. Beyond these results, he posed the problem of finding a formula for general $N(P_{n}\times P_{n})$ with noting the difficulty of solving this problem in [21], since determining the number of connected sets in an arbitrary graph is #P-complete [18].
Problem 1.1 ([21]).
Find a formula for the number of connected sets in the $(n\times n)$-lattice graph.
As a preliminary study to Problem 1.1, the authors of this paper have developed a recursive method in [16] to determine the numbers of connected sets in $(3\times n)$- and $(4\times n)$-lattice graphs, accompanied by the explicit enumeration formula for $N(K_{m}\times P_{n})$ (refer to Theorem 2.1) by employing the transfer matrix method, where $K_{m}$ is the complete graph of order $m$.
In this paper, we continue the study initiated in [16, 21] by further investigating the enumeration problem on connected sets in (cylindrical) lattice graphs. We first apply an approach in Section 2 similar to that in [16] to derive the enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$ (in Theorem 2.4), and provide detailed calculation results for the cases $m=4,5$ as examples in Section 3 (see Corollaries 3.1 and 3.2). Since these enumeration formulas are not in closed form, we further focus on establishing explicit lower bounds of $N(G\times P_{n})$ for any graph $G$ (see Theorem 1.2 below) by the transfer matrix method.
For any graph $G$, we call a maximal set in $G\times P_{n}$ consisting of all the vertices with the same second coordinate a column. Then any column induces a copy of $G$ in $G\times P_{n}$, and two columns $X,Y$ are said to be consecutive if there exist $(t,u)\in X$ and $(t,v)\in Y$ such that $uv\in E(P_{n})$. Let $N_{L}(G\times P_{n})$ be the number of connected sets $C$ of $G\times P_{n}$ that induce connected sets in $G\times P_{n}[X\cup Y]$ for every pair of consecutive columns $X,Y$ of $G\times P_{n}$ with $X\cap C,Y\cap C\neq\emptyset$. Then it is easy to see that $N_{L}(G\times P_{n})$ is a natural lower bound for $N(G\times P_{n})$. Moreover, $N_{L}(G\times P_{n})$ is a tight lower bound and its explicit enumeration formula is as follows.
Theorem 1.2.
Let $G$ be a connected graph of order $m$. Then
$N(G\times P_{n})\geq N_{L}(G\times P_{n})=n\cdot N(G)+\sum\limits_{k=2}^{n}(n-k+1)\cdot(\mathbf{1}^{T}\cdot A^{k-1}\mathbf{1}),$ (1.1)
where $A=(a_{ij})_{i,j\in\{1,2,...,2^{m}-1\}}$, $\phi$ is a fixed bijection from $\{1,2,...,2^{m}-1\}$ to $2^{V(G)}\setminus\{\emptyset\}$, $uv\in E(P_{n})$, and
$a_{ij}=\begin{cases}1,&\text{if}~\{(p,u):p\in\phi(i)\}\cup\{(q,v):q\in\phi(j)\}\in\mathcal{C}(G\times P_{n}),\\ 0,&\text{otherwise}.\end{cases}$
Moreover, the equality in (1.1) holds only when $G$ is a complete graph or $n\in\{1,2\}$.
The proof of Theorem 1.2 will be given in Section 4. By Theorem 1.2, we further provide in Section 4 an asymptotic analysis on the behavior of $N(P_{3}\times P_{n})$, $N(C_{4}\times P_{n})$, $N(C_{5}\times P_{n})$ and $N(K_{1,3}\times P_{n})$ with respect to parameters $c(P_{3}\times P_{n})$, $c(C_{4}\times P_{n})$, $c(C_{5}\times P_{n})$ and $c(K_{1,3}\times P_{n})$ as $n$ tends to infinity, where $K_{1,3}$ is the star graph of order four. Note that a similar analysis can also be done for other Cartesian product graphs $G\times P_{n}$, although the calculations become more time-consuming as $|V(G)|$ increases.
2 The enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$
In this section, we shall determine the enumeration formula for $N(C_{m}\times P_{n})$ with $m\leq 7$.
For $G=P_{m}\times P_{n}$ or $C_{m}\times P_{n}$, it is clear to see that $\mathcal{C}(G)\subseteq\mathcal{C}(K_{m}\times P_{n})$ as $G$ is a subgraph of $K_{m}\times P_{n}$. Since to determine the enumeration formula for $N(K_{m}\times P_{n})$ is rather simple (see Theorem 2.1), a natural approach here is to identify $\mathcal{C}(G)$ by removing some elements from $\mathcal{C}(K_{m}\times P_{n})$, in which any such element must contain a pair of vertices that are adjacent in $K_{m}\times P_{n}$ but not adjacent in $G$ (the reverse may not be true). This idea has been applied to derive the enumeration formula for $N(P_{m}\times P_{n})$ with $m\leq 4$ in [16]. In the following, we shall extend this approach to obtain similar results for $N(C_{m}\times P_{n})$ with $m\leq 7$ based on the background graph $K_{m}\times P_{n}$.
We first fix a labeling of $V(K_{m}\times P_{n})$ for all $m,n$, which naturally induces a labeling of $V(C_{m}\times P_{n})$ and $V(P_{m}\times P_{n})$, and will be used throughout Sections 2 and 3. Assume that $V(K_{m}\times P_{n})=\{v_{i,j}:i\in\{1,2,...,m\},j\in\{1,2,...,n\}\}$, where for any $i_{1},i_{2}\in\{1,2,...,m\}$ and $j_{1},j_{2}\in\{1,2,...,n\}$, $v_{i_{1},j_{1}}v_{i_{2},j_{1}}\in E(K_{m}\times P_{n})$, and $v_{i_{1},j_{1}}v_{i_{1},j_{2}}\in E(K_{m}\times P_{n})$ if and only if $|j_{1}-j_{2}|=1$. See an example in Figure 2. Then for each $j=1,2,\dots,n$, $I_{j}=\{v_{i,j}:i\in\{1,2,...,m\}\}$ is a column, and we call it the $j$-th column. For any $S_{j}\subseteq I_{j}$ with $j\geq 2$, let $S_{j-1}$ be the set of vertices in $I_{j-1}$ that are adjacent to the vertices in $S_{j}$. Then $|S_{j}|=|S_{j-1}|$ clearly holds.
[0.45]Refer to caption
Figure 2: The graph $K_{4}\times P_{n}$
Let $\mathcal{C}_{m,n}$ denote the subset of connected sets of $K_{m}\times P_{n}$ that contain at least one vertex from each column. Then it is easy to see that $N(K_{m}\times P_{n})=\sum\limits_{k=1}^{n}(n-k+1)\cdot|\mathcal{C}_{m,k}|.$ For any $i=1,2,\dots,m$, let $f_{m,n}^{i}$ denote the number of sets in $\mathcal{C}_{m,n}$ that contain a fixed set of $i$ vertices in the $n$-th column. Then as illustrated in Theorem 1.2 of [16], $|\mathcal{C}_{m,n}|=\sum\limits_{i=1}^{m}\binom{m}{i}f_{m,n}^{i}$, and $f_{m,n}^{i}$ can be derived by the recurrence formula that
$f_{m,n}^{i}=|\mathcal{C}_{m,n-1}|-\sum\limits_{j=1}^{m-i}\binom{m-i}{j}f_{m,n-1}^{j},$ (2.2)
with $f_{m,1}^{i}=1$ for all $i=1,2,\dots,m$. Hence, we present the counting formula for $N(K_{m}\times P_{n})$.
Theorem 2.1 ([16]).
For any positive integers $m$ and $n$,
$N(K_{m}\times P_{n})=\sum\limits_{k=1}^{n}(n-k+1)\cdot|\mathcal{C}_{m,k}|,$
where $|\mathcal{C}_{m,k}|=\begin{bmatrix}\binom{m}{1}&\binom{m}{2}&\cdots&\binom{m}{m}\end{bmatrix}\cdot T^{k-1}\cdot\mathbf{1}$, $T=(t_{ij})_{i,j\in\{1,2,...,m\}}$, and
$t_{ij}=\begin{cases}\binom{m}{j}-\binom{m-i}{j},&\text{when }j\leq m-i,\\ \binom{m}{j},&\text{when }j\geq m-i+1.\end{cases}$
By Theorem 2.1, we shall now focus on the set $\mathcal{C}(K_{m}\times P_{n})\setminus\mathcal{C}(C_{m}\times P_{n})$. Let $\mathcal{C}_{m,n}^{\prime}$ be the subset of $\mathcal{C}(K_{m}\times P_{n})\setminus\mathcal{C}(C_{m}\times P_{n})$, where each element of $\mathcal{C}_{m,n}^{\prime}$ contains at least one vertex from every column of $K_{m}\times P_{n}$. Clearly, $\mathcal{C}_{m,n}^{\prime}\subseteq\mathcal{C}_{m,n}$. For any nonempty set $X\subseteq V(K_{m}\times P_{n})$, let
$N(K_{m}\times P_{n};X)$ $=$ $|\{C\in\mathcal{C}_{m,n}:C\cap I_{i}=X\cap I_{i}~\text{whenever}~X\cap I_{i}\neq\emptyset\}|,$
$N^{\prime}(K_{m}\times P_{n};X)$ $=$ $|\{C\in\mathcal{C}_{m,n}^{\prime}:C\cap I_{i}=X\cap I_{i}~\text{whenever}~X\cap I_{i}\neq\emptyset\}|.$
Then $N(K_{m}\times P_{n};X)=f_{m,n}^{|X|}$ when $X\subseteq I_{n}$, and the following lemma is easy to observe.
Lemma 2.2.
For any positive integers $m$ and $n$,
$|\mathcal{C}_{m,n}^{\prime}|=\sum_{\emptyset\neq S_{n}\subseteq I_{n}}N^{\prime}(K_{m}\times P_{n};S_{n}),$ (2.3)
where
$N^{\prime}(K_{m}\times P_{n};S_{n})=\sum_{\emptyset\neq X\subseteq I_{n-1}\atop X\cap S_{n-1}\neq\emptyset}N^{\prime}(K_{m}\times P_{n};S_{n}\cup X)~\text{when}~n\geq 2.$ (2.4)
By (2.3) and (2.4), we are able to calculate the value of $|\mathcal{C}_{m,n}^{\prime}|$ when $n=1,2$. For the case $n\geq 3$, we shall deal with the following recursive lemma, which can be derived in a manner parallel to Lemma 3.2 in [16] and the proof is thus omitted. Here we denote by $k(G)$ the number of components in $G$ for any graph $G$.
Lemma 2.3.
For any $n\geq 3$ and $C\in\mathcal{C}_{m,n}^{\prime}$ with $C\cap I_{n}=S_{n}$ and $C\cap I_{n-1}=T_{n-1}$,
  1. (1)
    if some component of $C_{m}\times P_{n}[S_{n}]$ is also a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$, then $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1})=N(K_{m}\times P_{n-1};T_{n-1})=f_{m,n-1}^{|T_{n-1}|}$;
  2. (2)
    if each component of $C_{m}\times P_{n}[S_{n}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ and $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}])=k(C_{m}\times P_{n-1}[T_{n-1}])$, then $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1})=N^{\prime}(K_{m}\times P_{n-1};T_{n-1})$;
  3. (3)
    for the remaining case that each component of $C_{m}\times P_{n}[S_{n}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ and $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}])<k(C_{m}\times P_{n-1}[T_{n-1}])$, we have
    $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1})=\sum_{\emptyset\neq X\subseteq I_{n-2}\atop X\cap T_{n-2}\neq\emptyset}N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X).$ (2.5)
    Moreover,
    1. (i)
      if some component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ is also a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$, then $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X)=N(K_{m}\times P_{n-2};X)=f_{m,n-2}^{|X|}$;
    2. (ii)
      if each component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$ and $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X])=k(C_{m}\times P_{n-2}[X])$, then $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X)=N^{\prime}(K_{m}\times P_{n-2};X)$;
    3. (iii)
      otherwise, each component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$ and $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X])<k(C_{m}\times P_{n-2}[X])$. By the symmetry of $C_{m}$, assume that $R_{1},R_{2},\dots,R_{r}$ are the components of $C_{m}\times P_{n-2}[X]$, where $k_{s}<k_{t}$ when $s<t$ for all vertices $v_{k_{s},n-2}$ in $R_{s}$ and $v_{k_{t},n-2}$ in $R_{t}$. For each $i=1,2,\dots,r$, let $\psi(i)=\{j:R_{j}~\text{lies in the same component as}~R_{i}~\text{in}~C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]\}$, and let $l_{i}$ (or $\overline{l}_{i}$) be the smallest (or largest) integer $q$ such that $v_{q,n-2}$ belongs to the component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$ containing $R_{i}$. If $\psi(i)$ is a set of consecutive numbers for each $i=1,2,\dots,r$, then $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X)=N^{\prime}(K_{m}\times P_{n-1};X\cup\{v_{q,n-1}:l_{i}\leq q\leq\overline{l}_{i},i=1,\dots,r\})$.
Remark. Note that for Lemma 2.3 (3), the only exceptional case is characterized by the following: each component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}]$ is not a component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$, $k(C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X])<k(C_{m}\times P_{n-2}[X]),$ and there exist four components $R_{p},R_{i},R_{q},R_{j}$ in $C_{m}\times P_{n-2}[X]$ with $p<i<q<j$ such that $R_{p},R_{q}$ are in the same component of $C_{m}\times P_{n}[S_{n}\cup T_{n-1}\cup X]$ while $R_{i},R_{j}$ are in other components. An example is shown in Figure 3. For such cases, it is difficult to find a recurrence formula to calculate $N^{\prime}(K_{m}\times P_{n};S_{n}\cup T_{n-1}\cup X)$. However, for the case $m\leq 7$, we have $k(C_{m}\times P_{n-2}[X])\leq 3$ for any $X\subseteq I_{n-2}$, which implies that the situation described above does not arise.
[0.7]Refer to caption
Figure 3: An example of the case that Lemma 2.3 can not deal with, where $C$ is the set of solid vertices
Now it is clear to see that Lemmas 2.2 and 2.3 together provide a lower bound of $|{\cal C}_{m,n}^{\prime}|$ for all $m,n$, which can be calculated by multi-step recurrence formulas involving (2.2). Then by Theorem 2.1, we obtain a sharp upper bound for $N(C_{m}\times P_{n})$.
Theorem 2.4.
For any positive integers $m$ and $n$,
$N(C_{m}\times P_{n})\leq\sum\limits_{k=1}^{n}(|{\cal C}_{m,k}|-|{\cal C}_{m,k}^{\prime}|)\cdot(n-k+1),$
where the equality holds when $m\leq 7$.
3 Calculation results on $N(C_{4}\times P_{n})$ and $N(C_{5}\times P_{n})$
In this section, we shall present detailed calculation results for Theorem 2.4 when $m=4,5$.
In what follows, we abbreviate $N(K_{m}\times P_{n};\{x_{1},x_{2},...,x_{j}\})$ and $N^{\prime}(K_{m}\times P_{n};\mathcal{\{}x_{1},x_{2},...,x_{j}\})$ as $N(K_{m}\times P_{n};x_{1},x_{2},...,x_{j})$ and $N^{\prime}(K_{m}\times P_{n};x_{1},x_{2},...,x_{j})$, respectively.
We first focus on $N(C_{4}\times P_{n})$. For any positive integer $n$, assume that $v_{i+4,n}=v_{i,n}$ for all $i=1,2,3,4$, and let
  • $a_{n}=N^{\prime}(K_{4}\times P_{n};v_{i,n})$ for $i=1,2,3,4$;
  • $b_{n}^{1}=N^{\prime}(K_{4}\times P_{n};v_{i,n},v_{i+1,n})$ and $b_{n}^{2}=N^{\prime}(K_{4}\times P_{n};v_{i,n},v_{i+2,n})$ for $i=1,2,3,4$;
  • $c_{n}=N^{\prime}(K_{4}\times P_{n};v_{i,n},v_{i+1,n},v_{i+2,n})$ for $i=1,2,3,4$;
  • $d_{n}=N^{\prime}(K_{4}\times P_{n};I_{n})$;
  • $x_{n}=N^{\prime}(K_{4}\times P_{n};I_{n}\cup\{v_{1,n-1},v_{3,n-1}\})$ when $n\geq 2$.
Then the following corollary is a special case of Theorem 2.4.
Corollary 3.1.
For any positive integer $n$,
$N(C_{4}\times P_{n})=\sum\limits_{k=1}^{n}(|\mathcal{C}_{4,k}|-|\mathcal{C}_{4,k}^{\prime}|)\cdot(n-k+1),$
where
  1. (i)
    $|\mathcal{C}_{4,k}|=\begin{bmatrix}\binom{4}{1}&\binom{4}{2}&\binom{4}{3}&\binom{4}{4}\end{bmatrix}\cdot\left[\begin{matrix}1&3&3&1\\ 2&5&4&1\\ 3&6&4&1\\ 4&6&4&1\end{matrix}\right]^{k-1}\cdot\mathbf{1};$
  2. (ii)
    $|\mathcal{C}_{4,k}^{\prime}|=4a_{k}+4b_{k}^{1}+2b_{k}^{2}+4c_{k}+d_{k},$
    with $a_{1}=b_{1}^{1}=c_{1}=d_{1}=x_{2}=0$, $b_{1}^{2}=1$, and when $k\geq 2$,
    $\left\{\begin{aligned} a_{k}&=a_{k-1}+2b_{k-1}^{1}+b_{k-1}^{2}+3c_{k-1}+d_{k-1},\\ b_{k}^{1}&=2a_{k-1}+3b_{k-1}^{1}+2b_{k-1}^{2}+4c_{k-1}+d_{k-1},\\ b_{k}^{2}&=2f_{4,k-1}^{1}+4f_{4,k-1}^{2}+b_{k-1}^{2}+2c_{k-1}+2f_{4,k-1}^{3}+d_{k-1},\\ c_{k}&=3a_{k-1}+4b_{k-1}^{1}+b_{k-1}^{2}+x_{k}+4c_{k-1}+d_{k-1},\\ d_{k}&=4a_{k-1}+4b_{k-1}^{1}+2x_{k}+4c_{k-1}+d_{k-1},\\ x_{k}&=2a_{k-2}+4b_{k-2}^{1}+x_{k-1}+4c_{k-2}+d_{k-2}~\text{where}~k\geq 3.\end{aligned}\right$.
Proof.
By Theorems 2.1 and 2.4, we only need to verify (ii). Then due to (2.3), it suffices to verify the formulas for $a_{k},b_{k}^{1},b_{k}^{2},c_{k},d_{k}$ along with their respective initial values. The case when $k=1$ is obvious. Assume $k\geq 2$ in the following.
By (2.4) and Lemma 2.3 (2), we have
$a_{k}$ $=$ $N^{\prime}(K_{4}\times P_{k};v_{1,k})$
$=$ $N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1},v_{2,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1},v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1},v_{2,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1},v_{2,k-1},v_{4,k-1})$
$N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{1,k-1},v_{3,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k};\{v_{1,k}\}\cup I_{k-1})$
$=$ $N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k-1},v_{2,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1},v_{4,k-1})$
$N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{3,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};I_{k-1})$
$=$ $a_{k-1}+2b_{k-1}^{1}+b_{k-1}^{2}+3c_{k-1}+d_{k-1}.$
Similarly, we deduce that
$b_{k}^{1}$ $=$ $N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{2,k})$
$=$ $N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1},v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1},v_{3,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{3,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{4,k-1},v_{2,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};I_{k-1})$
$=$ $2a_{k-1}+3b_{k-1}^{1}+2b_{k-1}^{2}+4c_{k-1}+d_{k-1}.$
By (2.4) and Lemma 2.3 (1), (2), we have
$b_{k}^{2}$ $=$ $N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k})$
$=$ $N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{1,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{1,k-1},v_{2,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{1,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{2,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{3,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{1,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{1,k-1},v_{2,k-1},v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{2,k-1},v_{3,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{3,k-1},v_{4,k-1},v_{1,k-1})$
$+N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{3,k},v_{4,k-1},v_{1,k-1},v_{2,k-1})+N^{\prime}(K_{4}\times P_{k};\{v_{1,k},v_{3,k}\}\cup I_{k-1})$
$=$ $N(K_{4}\times P_{k-1};v_{1,k-1})+N(K_{4}\times P_{k-1};v_{3,k-1})$
$+N(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1})+N(K_{4}\times P_{k-1};v_{1,k-1},v_{4,k-1})$
$+N(K_{4}\times P_{k-1};v_{2,k-1},v_{3,k-1})+N(K_{4}\times P_{k-1};v_{3,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1},v_{3,k-1})$
$+N(K_{4}\times P_{k-1};v_{2,k-1},v_{3,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{3,k-1},v_{4,k-1},v_{1,k-1})$
$+N(K_{4}\times P_{k-1};v_{4,k-1},v_{1,k-1},v_{2,k-1})+N^{\prime}(K_{4}\times P_{k-1};I_{k-1})$
$=$ $2f_{4,k-1}^{1}+4f_{4,k-1}^{2}+b_{k-1}^{2}+2f_{4,k-1}^{3}+2c_{k-1}+d_{k-1},$
where $f_{4,k-1}^{i}$ with $i=1,2,3$ can be calculated recursively by (2.2).
Next, by (2.4) and Lemma 2.3 (2), (3), we have
$c_{k}$ $=$ $N^{\prime}(K_{4}\times P_{k};v_{1,k},v_{2,k},v_{3,k})$
$=$ $N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1})+x_{n}+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1},v_{3,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{3,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1},v_{3,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{2,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};v_{2,k-1},v_{3,k-1},v_{4,k-1})$
$+N^{\prime}(K_{4}\times P_{k-1};v_{1,k-1},v_{3,k-1},v_{4,k-1})+N^{\prime}(K_{4}\times P_{k-1};I_{k-1})$
$=$ $3a_{k-1}+4b_{k-1}^{1}+b_{k-1}^{2}+x_{k}+4c_{k-1}+d_{k-1},$
and
$d_{k}$ $=$ $N^{\prime}(K_{4}\times P_{k};I_{k})=4a_{k-1}+4b_{k-1}^{1}+2x_{k}+4c_{k-1}+d_{k-1}.$
Then it remains to verify the recurrence formula for $x_{k}$. Obviously, $x_{2}=0$. For $k\geq 3$, by Lemma 2.3 (3), we have
$x_{k}$ $=$ $N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1}\})$
$=$ $N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2}\})+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{3,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{2,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{4,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{2,k-2},v_{3,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{3,k-2},v_{4,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{3,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{2,k-2},v_{3,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{2,k-2},v_{4,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{3,k-2},v_{4,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{2,k-2},v_{3,k-2},v_{4,k-2}\})$
$+N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1}\}\cup I_{k-2}),$
where
$N^{\prime}(K_{4}\times P_{k};I_{k}\cup\{v_{1,k-1},v_{3,k-1},v_{1,k-2},v_{3,k-2}\})$
$=$ $N^{\prime}(K_{4}\times P_{k-1};I_{k-1}\cup\{v_{1,k-2},v_{3,k-2}\})$
$=$ $x_{k-1}.$
Thus,
$x_{k}=2a_{k-2}+4b_{k-2}^{1}+x_{k-1}+4c_{k-2}+d_{k-2}.$
The proof is complete. ∎
Similarly, we can establish the recurrence formulas for $N(C_{5}\times P_{n})$. For any positive integer $n$, assume that $v_{i+5,n}=v_{i,n}$ for all $i=1,2,3,4,5$, and let
  • $a_{n}=N^{\prime}(K_{5}\times P_{n};v_{i,n})$ for $i=1,2,3,4,5$;
  • $b_{n}^{1}=N^{\prime}(K_{5}\times P_{n};v_{i,n},v_{i+1,n})$ and $b_{n}^{2}=N^{\prime}(K_{5}\times P_{n};v_{i,n},v_{i+2,n})$ for $i=1,2,3,4,5$;
  • $c_{n}^{1}=N^{\prime}(K_{5}\times P_{n};v_{i,n},v_{i+1,n},v_{i+2,n})$ and $c_{n}^{2}=N^{\prime}(K_{5}\times P_{n};v_{i,n},v_{i+2,n},v_{i+3,n})$ for $i=1,2,3,4,5$;
  • $d_{n}=N^{\prime}(K_{5}\times P_{n};v_{i,n},v_{i+1,n},v_{i+2,n},v_{i+3,n})$ for $i=1,2,3,4,5$;
  • $g_{n}=N^{\prime}(K_{5}\times P_{n};I_{n})$;
  • $x_{n}=N^{\prime}(K_{5}\times P_{n};v_{1,n},v_{2,n},v_{3,n},v_{1,n-1},v_{3,n-1})$ when $n\geq 2$;
  • $y_{n}=N^{\prime}(K_{5}\times P_{n};v_{1,n},v_{2,n},v_{3,n},v_{1,n-1},v_{3,n-1},v_{4,n-1})\\ \hskip-227.62204pt=N^{\prime}(K_{5}\times P_{n};v_{1,n},v_{2,n},v_{3,n},v_{1,n-1},v_{3,n-1},v_{5,n-1})$ when $n\geq 2$.
Then the next corollary is direct.
Corollary 3.2.
For any positive integer $n$,
$N(C_{5}\times P_{n})=\sum\limits_{k=1}^{n}(|\mathcal{C}_{5,k}|-|\mathcal{C}_{5,k}^{\prime}|)\cdot(n-k+1),$
where
  1. (i)
    $|\mathcal{C}_{5,k}|=\begin{bmatrix}\binom{5}{1}&\binom{5}{2}&\binom{5}{3}&\binom{5}{4}&\binom{5}{5}\end{bmatrix}\cdot\left[\begin{matrix}1&4&6&4&1\\ 2&7&9&5&1\\ 3&9&10&5&1\\ 4&10&10&5&1\\ 5&10&10&5&1\end{matrix}\right]^{k-1}\cdot\mathbf{1};$
  2. (ii)
    $|\mathcal{C}_{5,k}^{\prime}|=5(a_{k}+b_{k}^{1}+b_{k}^{2}+c_{k}^{1}+c_{k}^{2}+d_{k})+g_{k},$
    with $a_{1}=b_{1}^{1}=c_{1}^{1}=d_{1}=g_{1}=x_{2}=y_{2}=0,b_{1}^{2}=c_{1}^{2}=1$, and when $k\geq 2$,
    $\left\{\begin{aligned} a_{k}&=a_{k-1}+2b_{k-1}^{1}+2b_{k-1}^{2}+3c_{k-1}^{1}+3c_{k-1}^{2}+4d_{k-1}+g_{k-1},\\ b_{k}^{1}&=2a_{k-1}+3b_{k-1}^{1}+4b_{k-1}^{2}+4c_{k-1}^{1}+5c_{k-1}^{2}+5d_{k-1}+g_{k-1},\\ b_{k}^{2}&=2f_{5,k-1}^{1}+6f_{5,k-1}^{2}+b_{k-1}^{2}+6f_{5,k-1}^{3}+c_{k-1}^{1}+2c_{k-1}^{2}+2f_{5,k-1}^{4}+3d_{k-1}+g_{k-1},\\ c_{k}^{1}&=3a_{k-1}+4b_{k-1}^{1}+x_{k}+4b_{k-1}^{2}+5c_{k-1}^{1}+3c_{k-1}^{2}+2y_{k}+5d_{k-1}+g_{k-1},\\ c_{k}^{2}&=3f_{5,k-1}^{1}+7f_{5,k-1}^{2}+2b_{k-1}^{2}+5f_{5,k-1}^{3}+2c_{k-1}^{1}+3c_{k-1}^{2}+f_{5,k-1}^{4}+4d_{k-1}+g_{k-1},\\ d_{k}&=4a_{k-1}+5b_{k-1}^{1}+2b_{k-1}^{2}+3x_{k}+5c_{k-1}^{1}+4y_{k}+c_{k-1}^{2}+5d_{k-1}+g_{k-1},\\ g_{k}&=5a_{k-1}+5b_{k-1}^{1}+5x_{k}+5y_{k}+5c_{k-1}^{1}+5d_{k-1}+g_{k-1},\\ x_{k}&=2a_{k-2}+4b_{k-2}^{1}+x_{k-1}+2b_{k-2}^{2}+5c_{k-2}^{1}+2c_{k-2}^{2}+2y_{k-1}+5d_{k-2}+g_{k-2}~\text{where}~k\geq 3,\\ y_{k}&=3a_{k-2}+5b_{k-2}^{1}+2x_{k-1}+2b_{k-2}^{2}+5c_{k-2}^{1}+2c_{k-2}^{2}+3y_{k-1}+5d_{k-2}+g_{k-2}~\text{where}~k\geq 3.\end{aligned}\right$.
4 Asymptotic analysis
In this section, we prove Theorem 1.2 and perform an asymptotic analysis on $N(P_{3}\times P_{n})$, $N(C_{4}\times P_{n})$, $N(C_{5}\times P_{n})$ and $N(K_{1,3}\times P_{n})$.
Proof of Theorem 1.2. We begin by proving (1.1). For each $k=1,2,\dots,n$, assume that $V(P_{k})=\{v_{1},v_{2},\dots,v_{k}\}$, where $v_{i}v_{i+1}\in E(P_{k})$ for all $i\in\{1,2,\dots,k-1\}$. Then the $j$-th column $I_{j}$ of $G\times P_{k}$ is clearly $\{(x,v_{j}):x\in V(G)\}$ for all $j\in\{1,2,\dots,k\}$.
Let $\mathcal{C}_{L}(k)$ be the set of sets $C$ in $\mathcal{C}(G\times P_{k})$ such that $C\cap I_{j}\neq\emptyset$ for each $j\in\{1,2,...,k\}$ and $(C\cap I_{j})\cup(C\cap I_{j+1})\in\mathcal{C}(G\times P_{k})$ for each $j\in\{1,2,...,k-1\}$. Then it suffices to show that $\mathbf{1}^{T}\cdot A^{k-1}\cdot\mathbf{1}=|\mathcal{C}_{L}(k)|$ for all $k\geq 2$.
In fact, we shall show by induction that for each $(i,j)$-entry in $A^{k-1}$, it is the number of sets $C$ in $\mathcal{C}_{L}(k)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{k}):q\in\phi(j)\}$. Then (1.1) is proven by the definition of $\phi$.
When $k=2$, the matrix $A$ clearly satisfies our requirements. Suppose the statement is true for $k=t$, where $t\geq 2$. Let $l_{i,r}$ be any $(i,r)$-entry in $A^{t-1}$. Then $l_{i,r}$ is the number of sets in $\mathcal{C}_{L}(t)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{t}):q\in\phi(r)\}$. Now we consider the case for $k=t+1$. Each $(i,j)$-entry in the matrix $A^{t}$ is $\sum\limits_{r=1}^{2^{m}-1}l_{i,r}a_{r,j}$, where $a_{r,j}=1$ if and only if $\{(p,v_{t}):p\in\phi(r)\}\cup\{(q,v_{t+1}):q\in\phi(j)\}\in\mathcal{C}(G\times P_{t+1}).$ Consequently, $\sum\limits_{r=1}^{2^{m}-1}l_{i,r}a_{r,j}$ is the number of sets in $\mathcal{C}_{L}(t+1)$ containing $\{(p,v_{1}):p\in\phi(i)\}\cup\{(q,v_{t+1}):q\in\phi(j)\}$. Hence the matrix $A^{t}$ satisfies the requirements in the statement and (1.1) is proven.
Next, we show that (1.1) is sharp only when $G$ is complete or $n\in\{1,2\}$. It is clear that the equality in (1.1) holds when $G$ is complete or $n\in\{1,2\}$. Now, assume that the equality in (1.1) holds for a connected graph $G$ that is not complete and $n\geq 3$. Then there exist nonadjacent vertices $u$ and $v$ in $G$ connected by a path $ux_{1}x_{2}\dots x_{t}v$ in $G$. Let
$C=\{(u,v_{1}),(v,v_{1}),(u,v_{2}),(v,v_{2}),(u,v_{3}),(x_{1},v_{3}),(x_{2},v_{3}),\dots,(x_{t},v_{3}),(v,v_{3})\}.$
Since $n\geq 3$, it can be easily verified that $C\in\mathcal{C}(G\times P_{n})$ and $C\cap I_{1},C\cap I_{2}\neq\emptyset$. However, $C\cap(I_{1}\cup I_{2})\notin\mathcal{C}(G\times P_{n})$. Hence, $N_{L}(G\times P_{n})<N(G\times P_{n})$, a contradiction. As a result, (1.1) is sharp only when $G$ is complete or $n\in\{1,2\}$. $\Box$
Note that the transfer matrix $A$ given in Theorem 1.2 is symmetric and makes the calculation of an estimate of $N(G\times P_{n})$ for any graph $G$ possible. In particular, this kind of estimate works well for (cylindrical) lattice graphs. See several examples below.
Example 4.1.
  1. (i)
    For $C_{4}\times P_{n}$, we have $A_{1}$ as the matrix $A$ in Theorem 1.2 below:
[…]
    Since $A_{1}$ is symmetric with nonnegative entries, the limit
    $\lim_{n\to\infty}\left(\mathbf{1}^{T}A_{1}^{n-1}\mathbf{1}\right)^{1/n}$
    exists by the Gelfand theorem, and is equal to the largest eigenvalue $\lambda_{A_{1}}$ of $A_{1}$. Using Mathematica, we find that $\lambda_{A_{1}}\approx 10.5318$. Hence, we obtain the following bound for $c(C_{4}\times P_{n})$:
    $c(C_{4}\times P_{n})>\lambda_{A_{1}}^{1/4}\approx 1.8014.$
  2. (ii)
    Similarly for $C_{5}\times P_{n}$, we have $\lambda_{A_{2}}\approx 18.2600$ with $A_{2}$ as the matrix $A$ in Theorem 1.2. Thus
    $c(C_{5}\times P_{n})>\lambda_{A_{2}}^{\frac{1}{5}}\approx 1.7877.$
  3. (iii)
    In fact, both bounds in (i) and (ii) provide a good estimate, as shown in Figure 4(a) and Table 1 that illustrate the asymptotic behavior of $c(C_{4}\times P_{n})$ and $c_{1}=(N_{L}(C_{4}\times P_{n}))^{1/4n}=\sqrt[4n]{13n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{1}^{k-1}\mathbf{1})\cdot(n-k+1)}$, and Figure 4(b) and Table 2 that show the asymptotic behavior of $c(C_{5}\times P_{n})$ and $c_{2}=(N_{L}(C_{5}\times P_{n}))^{1/5n}=\sqrt[5n]{21n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{2}^{k-1}\mathbf{1})\cdot(n-k+1)}$, when $1\leq n\leq 800$.
    (a) The comparison of $c(C_{4}\times P_{n})$ and $c_{1}$
    (b) The comparison of $c(C_{5}\times P_{n})$ and $c_{2}$
    Figure 4: Asymptotic behavior of $c$, $c_{1}$ and $c_{2}$
    Table 1: The values of $c(C_{4}\times P_{n})$ and $c_{1}$ for $1\leq n\leq 11$.
    $n$ 1 2 3 4 5 6 7 8 9 10 11
    $c(C_{4}\times P_{n})$ 1.8988 1.8960 1.8796 1.8690 1.8624 1.8580 1.8548 1.8525 1.8506 1.8492 1.8480
    $c_{1}$ 1.8988 1.8960 1.8706 1.8534 1.8430 1.8360 1.8310 1.8273 1.8244 1.8221 1.8202
    Table 2: The values of $c(C_{5}\times P_{n})$ and $c_{2}$ for $1\leq n\leq 11$.
    $n$ 1 2 3 4 5 6 7 8 9 10 11
    $c(C_{5}\times P_{n})$ 1.8384 1.8628 1.8687 1.8687 1.8683 1.8681 1.8679 1.8678 1.8677 1.8676 1.8676
    $c_{2}$ 1.8384 1.8628 1.8466 1.8308 1.8224 1.8165 1.8124 1.8093 1.8069 1.8049 1.8034
Example 4.2.
  1. (i)
    For $P_{3}\times P_{n}$, we have $A_{3}$ as the matrix $A$ in Theorem 1.2 below:
    $A_{3}=\begin{pmatrix}1&0&0&1&0&0&1\\ 0&1&0&1&1&0&1\\ 0&0&1&0&1&0&1\\ 1&1&0&1&1&0&1\\ 0&1&1&1&1&0&1\\ 0&0&0&0&0&0&1\\ 1&1&1&1&1&1&1\end{pmatrix}$.
    Then by using Mathematica, we have $\lambda_{A_{3}}\approx 4.6524$ and the following bound for $c(P_{3}\times P_{n})$:
    $c(P_{3}\times P_{n})>\lambda_{A_{3}}^{1/3}\approx 1.6694.$
  2. (ii)
    For $K_{1,3}\times P_{n}$, we have $\lambda_{A_{4}}\approx 8.9322$ with $A_{4}$ as the matrix $A$ in Theorem 1.2. Thus,
    $c(K_{1,3}\times P_{n})>\lambda_{A_{4}}^{1/4}\approx 1.7288.$
  3. (iii)
    In fact, both bounds in (i) and (ii) provide a good estimate, as shown in Figure 5(a) and Table 3 that illustrate the asymptotic behavior of $c(P_{3}\times P_{n})$ and $c_{3}=(N_{L}(P_{3}\times P_{n}))^{1/3n}=\sqrt[3n]{6n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{3}^{k-1}\mathbf{1})\cdot(n-k+1)}$, and Figure 5(b) and Table 4 show the asymptotic behavior of $c(K_{1,3}\times P_{n})$ and $c_{4}=(N_{L}(K_{1,3}\times P_{n}))^{1/4n}=\sqrt[4n]{11n+\sum\limits_{k=2}^{n}(\mathbf{1}^{T}A_{4}^{k-1}\mathbf{1})\cdot(n-k+1)}$, when $1\leq n\leq 800$.
    (a) The comparison of $c(P_{3}\times P_{n})$ and $c_{3}$
    (b) The comparison of $c(K_{1,3}\times P_{n})$ and $c_{4}$
    Figure 5: Asymptotic behavior of $c$, $c_{3}$ and $c_{4}$
Table 3: The values of $c(P_{3}\times P_{n})$ and $c_{3}$ for $1\leq n\leq 11$.
$n$ 1 2 3 4 5 6 7 8 9 10 11
$c(P_{3}\times P_{n})$ 1.8171 1.8493 1.8190 1.7960 1.7804 1.7697 1.7620 1.7563 1.7518 1.7482 1.7453
$c_{3}$ 1.8171 1.8493 1.8095 1.7772 1.7560 1.7414 1.7310 1.7231 1.7171 1.7123 1.7083
Table 4: The values of $c(K_{1,3}\times P_{n})$ and $c_{4}$ for $1\leq n\leq 11$.
$n$ 1 2 3 4 5 6 7 8 9 10 11
$c(K_{1,3}\times P_{n})$ 1.8212 1.8322 1.8224 1.8148 1.8098 1.8065 1.8042 1.8025 1.8011 1.8001 1.7992
$c_{4}$ 1.8212 1.8322 1.8075 1.7875 1.7757 1.7677 1.7621 1.7579 1.7547 1.7521 1.7499
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 12171402).
Statements and Declarations
  • Competing Interests: The authors have no relevant financial or non-financial interests to disclose.
  • Data availability: No data were used in this study.